OpenStudy (anonymous):

Given the function y=0.1x^3-0.2x^2+0.3x+1 Find the value of x when y=60.4 without any graphs.

OpenStudy (campbell_st):

well you have \[60.4 = 0.1x^3 -0.2x^2 + 0.3x + 1\] which becomes \[0 = 0.1x^2 - 0.2x^2 + 0.3x - 59.4\] multiply every term by 10 \[0 = x^3 - 2x^2 + 3x - 594\] I'd use a bit of guess and check x = 10 1000 - 200 + 30 - 594 = 236 x = 8 gives 8^3 -2*8^2 +3*8 - 594 = -186 so the zero is between 10 and 8 try x = 9 9^3 - 2x^2 +3*8 - 594 = 0 so the solution is x = 9

OpenStudy (anonymous):

haha, yeah. I was trying to avoid that. I found that there is a cubic formula but thanks for the help anyway. I will give you a medal. :)